Please Help
Which of the following ordered pairs are solutions of 4x-3y=27? select all that apply.
A.(-5,3)
B.(3,6)
C.(12,15 3/4)
D.(-1,7 3/4)
E.(0,6 3/4)
F.(9 3/4, 0)

Answer by funmath(2873)

Please Help
Which of the following ordered pairs are solutions of 4x-3y=27? select all that apply.
A.(-5,3)
B.(3,6)
C.(12,15 3/4)
D.(-1,7 3/4)
E.(0,6 3/4)
F.(9 3/4, 0)
:
For A, let x=-5 and y=3, if the equation is equal, it is a solution.
4(-5)-3(3)=27
-20-9=27
-29=27 No, this one is not a solution.
:
For B, let x=3 and y=6
4(3)-3(6)=27
12-18=27
-6=27 No, this one is not a solution.
:
For C, let x=12 and y=15 3/4=63/4
4(-5)-3(63/4)=27
-20-189/4=27
-80/4-189/4=27
-269/4=27 No.
:
For D, let x=-1 and y=7 3/4=31/4
4(-1)-3(31/4)=27
-4-93/4=27
-16/4-93/4=27
-109/4=27 no.
:
For E, let x=0 and y=6 3/4=27/4
4(0)-3(27/4)=27
0-81/4=27
-81/4=27 no.
:
For F, let x=9 3/4=31/4 and y=0
4(31/4)-3(0)=27
31-0=27
31=27 no.
:
Are you sure you typed everything correctly, because something should have worked.
Hppy Calculating!!!